OpenStudy (anonymous):

Select the proper equation in point-slope form for the given point and slope. Point: (7, 4); Slope: 6 y – 7 = 6(x + 4) y – 4 = 6(x – 7) y – 4 = 6(x + 7) y + 4 = 6(x – 7)

OpenStudy (yttrium):

Do you still remember point slope-form. \[y-y _{1} = m(x-x _{2})\] Just plug in the values.

OpenStudy (anonymous):

Kind of? I'm kind of brain dead- It's one am where ia m

OpenStudy (yttrium):

y1 = 4 x1 = 7 m = 6 Then what should be the answer?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

I'm still confused @Yttrium

OpenStudy (yttrium):

With the form I've mention upward^ and thru the givens. We can solve for its equation. y-y1 = m (x-x1) right? hence, y - 4 = 6 (x-7) that's it. :)

OpenStudy (anonymous):

y-4=6 (x+7) ?????

OpenStudy (anonymous):

ooooooooooh its (x-7) thats what confused me

OpenStudy (anonymous):

thanks a bunch c: @Yttrium

OpenStudy (yttrium):

No problem anyway.